I need help factoring 28-11t+t^2. I would be grateful for the help.
Answer by Theo(2978)

equation is:

t^2 - 11t + 28

4*7 = 28 and 4+7 = 11 so that's a good candidate.

make your factors (x-7) * (x-4) and you should be ok.

multiply the factors out to double check.

(x-7) * (x-4) =

x^2 - 4x - 7x + 28

combine like terms to get:

x^2 - 11x + 28

looks ok.

should be your answer.

I used x instead of t, but the concept is the same.

replace my x with your t and you'll be good to go.